# For Better or For Worse? Learning Minimum Variance Features With Label Augmentation

### v2

@@ Abstract
Data augmentation has been pivotal in successfully training deep learning models on classification tasks over the past decade. An important subclass of data augmentation techniques - which includes both label smoothing and Mixup - involves modifying not only the input data but also the input label during model training. In this work, we analyze the role played by the label augmentation aspect of such methods. We first prove that linear models on binary classification data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features. We then use our techniques to show that even for nonlinear models and general data distributions, the label smoothing and Mixup losses are lower bounded by a function of the model output variance. An important consequence of our results is negative: label smoothing and Mixup can be less robust to spurious correlations in the data. We verify that our theory reflects practice via experiments on image classification benchmarks modified to have spurious correlations.
@@ 1 Introduction
Although the connection between the label smoothing and Mixup losses has been noted in both theory and practice (Carratino et al., 2020), there has not been (to the best of our knowledge) a unifying theoretical perspective on why the models trained using these losses exhibit similar behavior. In this paper, we aim to provide such a perspective by extending the ideas from previous feature learning analyses to the aspect shared in common between both Mixup and label smoothing: label augmentation. Our analysis allows us to identify a clear failure mode shared by both Mixup and label smoothing - that they can ignore higher variance features - which we also verify empirically.
@@ 1.1 Main Contributions
Our main contributions are three-fold:
  1. 1.
    We introduce a simple class of data distributions in which there exists a low variance feature and a high variance feature, and show that optimizing the label smoothing and Mixup losses for a linear model for such distributions leads to learning only the low variance feature. On the other hand, if the high variance feature satisfies a greater separability condition relative to the low variance feature, optimizing the empirical risk with weight decay successfully learns both features in the data. Our label smoothing and Mixup results are stated under general conditions on both the data and the respective loss hyperparameters; our Mixup results apply to any mixing distribution that does not coincide with ERM, which greatly generalizes the settings considered by prior work on Mixup feature learning (Chidambaram et al., 2023; Zou et al., 2023).
  2. 2.
    We show that our techniques extend to the analysis of general model classes, and prove lower bounds for the label smoothing and Mixup losses in terms of the model output variance. These results suggest that optimizing these losses necessarily decreases model output variance, which is not necessarily the case for training with weight decay.
  3. 3.
    We empirically verify our theoretical insights by considering image classification benchmarks modified to (1) have adversarial spurious correlations in the training data and (2) more natural spurious correlations in the form of class-correlated backgrounds. In both cases, we observe that label smoothing and Mixup fail to generalize for the full range of hyperparameters used in practice, while ERM with weight decay does generalize.
@@ 1.2 Related Work
Label Smoothing. Since being introduced by Szegedy et al. (2015), label smoothing continues to be used to train SOTA vision (Wortsman et al., 2022; Liu et al., 2022), translation (Vaswani et al., 2017; Team et al., 2022), and multi-modal models (Yu et al., 2022). Attempts to understand when and why label smoothing is effective can be traced back to the influential empirical analyses of Pereyra et al. (2017) and Müller et al. (2020), which respectively relate label smoothing to entropy regularization and show that it can lead to more closely clustered learned representations, with the latter work further showing that these clustered representations can be responsible for decreased distillation performance.
On the theoretical side, Lukasik et al. (2020) study the relationship between label smoothing and loss correction techniques used to handle label noise, and show that label smoothing can be effective for mitigating label noise. Liu (2021) extends this line of work by analyzing how label smoothing can outperform loss correction in the context of memorization of noisy labels. Wei et al. (2022) further show that negative label smoothing can outperform traditional label smoothing in the presence of label noise. Xu et al. (2020) provide an alternative theoretical perspective, showing that label smoothing can improve convergence of stochastic gradient descent. However, to the best of our knowledge, there are no preexisting theoretical works attempting to understand how label smoothing affects feature learning.
@@
On the other hand, Guo et al. (2019) and Chidambaram et al. (2021) show that the regularization terms introduced by the Mixup loss can also have a negative impact due to the fact that Mixup-augmented points may coincide with existing training data points, leading to models that fail to minimize the original risk. Mixup has also been studied from the perspective of calibration (Thulasidasan et al., 2019), with theoretical results (Zhang et al., 2021; Chidambaram & Ge, 2023) showing that Mixup training can prevent miscalibration issues arising from ERM (even when post-training calibration methods are used with ERM).
Recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM. Most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al. (2023) and Zou et al. (2023) both showing that Mixup training can learn multiple features in a generative, multi-view (Allen-Zhu & Li, 2021) data model despite ERM failing to do so. We provide a more detailed comparison to these works when we discuss our main results in Section 3.
@@ 2 Preliminaries
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(X_{1:2},Y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h(\lambda,g,X_{1:2},Y_{1:2})\right].$ (2.8)
@@ 3.1 Linear Binary Classification
We begin first with the binary setting, in which we will consider a data model where a subset of the input dimensions correspond to a low variance feature and the complementary dimensions correspond to a high variance feature that is more separable. This setup is intended to be a simplistic model of spurious correlations that appear in image data; one could think of the low variance feature as a fixed background while the high variance feature corresponds to the actual object category of interest.
The goal in such cases is to ensure that we do not only learn the low variance feature, so that generalization is possible even under a shift of this feature (for example, recognizing the same object on different backgrounds). We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to a model that has only learned the low variance feature, whereas minimizing the binary cross-entropy with non-zero weight decay requires learning the high variance feature due to its greater separability. The overarching data model for this subsection is as follows.
@@
For convenience, for a vector $v\in\mathbb{R}^{d}$, we will use $v\in\mathcal{L}$ to mean that only the $\mathcal{L}$ dimensions of $v$ are non-zero. Definition 3.1 does not actually make any strong assumptions on $\pi$ yet; this is because the weight decay and label augmentation results have different, non-conflicting assumptions that we codify separately. For the former, we need the data to be more separable in the $\mathcal{H}$ dimensions than in the $\mathcal{L}$ dimensions, in the following sense.
@@
Assumption 3.2 is strong, but it actually does not imply linear separability, only that the $\mathcal{H}$ dimensions are in a sense better than the $\mathcal{L}$ dimensions for classification. For a simple example to provide intuition for the assumption, one can consider a 2-D distribution in which the $x$ coordinates for both classes are concentrated close to 0, but the $y$ coordinates are either large negative or large positive values depending on the class.
@@
On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$. We will show that for both label smoothing and Mixup, the optimal solutions have arbitrarily small correlation with the dimensions in $\mathcal{H}$ as the distribution of $YX$ concentrates. For these results, we require the following assumptions (but no longer need Assumption 3.2).
@@
We assume that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\Sigma_{YX,\mathcal{H}}\succ\rho\mathrm{Id}$ for some $\rho>0$. Here $\Sigma_{YX,\mathcal{H}}$ denotes the covariance matrix of $\Sigma_{YX}$ restricted to those rows and columns in $\mathcal{H}$.
The first part of Assumption 3.5 just ensures it is possible to obtain a good solution using the dimensions in $\mathcal{L}$ while the second part codifies the idea of $\mathcal{H}$ being a high (at least non-zero) variance feature. Observe that we have made no direct separability assumptions on the data, although it is true that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ the class-conditional supports are guaranteed to be linearly separable in $\mathcal{L}$. We now state both results.
@@
Proof Sketch. The proof follows the same outline as the label smoothing result, but there are now additional difficulties due to mixing inputs. In particular, we need to condition on $\lambda$ and then show strong convexity of the conditional loss in terms of $w^{\top}Z_{\lambda}$. We can then get a lower bound on the overall loss as an expectation with respect to $\lambda$ of this conditional loss. However, it is no longer immediate that there exists a stationary point $w\in\mathcal{L}$ that minimizes this lower bound. We prove the existence of such a stationary point by considering limiting behavior of the gradient of the loss as $w_{\mathcal{L}}$ tends to $\infty$ or $\infty$- ∞ in each component. After establishing a stationary point, we can show as with the label smoothing result that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ we can get a solution with $w\in\mathcal{L}$ that achieves the lower bound while solutions with $w_{\mathcal{H}}$ non-zero are bounded away from the lower bound.
@@ 3.2 General Multi-Class Classification
$\ell_{\mathrm{LS},\alpha}(g)\geq\mathrm{OPT}_{\mathrm{LS},\alpha}+C\sum_{y=1}^{k}\pi_{Y}(y)\mathrm{Var}\left(g(X)\mid Y=y\right).$ (3.1)
@@
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)\geq\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}+C\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda}\left[\mathrm{Var}\left(g(Z_{\lambda})\mid y_{1},y_{2},\lambda\right)\right].$ (3.2)
@@
Additionally, our results can be viewed as generalizing the observations made by Chidambaram et al. (2023), which were that Mixup can learn multiple features in the data when doing so decreases the variance of the learned predictor. In our case, we directly show that Mixup will have much larger correlation with the low variance feature in the data as opposed to the high variance feature. Our results also do not contradict the observations of Zou et al. (2023), which were that Mixup can learn both a “common” feature and a “rare” feature in the data; in their setup the common/rare features are fixed (zero variance) per class and concatenated with high variance noise.
@@ 4 Experiments
We now verify the predictions of our theoretical results in Section 3 across both binary and multi-class experimental setups. All experiments in this section were conducted on a single P100 GPU using PyTorch (Paszke et al., 2019) for model implementation. All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
4.1 Spurious Pixels in Binary Classification

### v3

@@ Abstract
Data augmentation has been pivotal in successfully training deep learning models on classification tasks over the past decade. An important subclass of data augmentation techniques - which includes both label smoothing and Mixup - involves modifying not only the input data but also the input label during model training. In this work, we analyze the role played by the label augmentation aspect of such methods. We first prove that linear models on binary classification data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features. We then use our techniques to show that even for nonlinear models and general data distributions, the label smoothing and Mixup losses are lower bounded by a function of the model output variance. Lastly, we demonstrate empirically that this aspect of label smoothing and Mixup can be a positive and a negative. On the one hand, we show that the strong performance of label smoothing and Mixup on image classification benchmarks is correlated with learning low variance hidden representations. On the other hand, we show that Mixup and label smoothing can be more susceptible to low variance spurious correlations in the training data.
@@ 1 Introduction
Although the connection between the label smoothing and Mixup losses has been noted in both theory and practice (Carratino et al., 2020), there has not been (to the best of our knowledge) a unifying theoretical perspective on why the models trained using these losses exhibit similar behavior. In this paper, we aim to provide such a perspective by extending the ideas from previous feature learning analyses to the aspect shared in common between both Mixup and label smoothing: label augmentation. Our analysis shows that both Mixup and label smoothing hone in on low variance features in the data, and we demonstrate empirically that this phenomenon occurs in both synthetic settings with spuriously correlated features as well as in the training of standard deep learning models for image classification benchmarks.
@@ 1.1 Main Contributions
The main message of our paper can be summarized as:
In data distributions where there are low variance and high variance latent features, both label smoothing and Mixup incentivize learning models that correlate more strongly with the low variance features.
We prove this message concretely in Section 3.1, where we characterize linear optimizers of the label smoothing and Mixup losses on data in which some dimensions have lower variance than others. In Section 3.2, we also prove weaker analogues of our linear results in the context of general models and multi-class distributions; namely, we show that optimizing the label smoothing and Mixup losses requires decreasing model output variance in a way that is not necessarily true for empirical risk minimization (ERM) combined with weight decay.
We verify our theory empirically in Section 4, where we show that label smoothing and Mixup do indeed learn lower variance hidden representations (corroborating ideas from the literature on neural collapse (Papyan et al., 2020)) and that this lower variance property correlates with better generalization performance. We hypothesize that, for standard benchmarks, there exist low variance latent features that generalize well as opposed to high-variance, noisy features, and we expect that directly investigating such features would be a fruitful line of future work. However, we also point out that our key message does not directly imply better generalization performance – we show via synthetic experiments in Section 4.2 that it is possible for the low variance features to actually be spuriously correlated with the targets (e.g. fixed backgrounds).
@@ 1.2 Related Work
Label Smoothing. Since being introduced by Szegedy et al. (2015), label smoothing continues to be used to train SOTA vision (Wortsman et al., 2022; Liu et al., 2022), translation (Vaswani et al., 2017; Team et al., 2022), and multi-modal models (Yu et al., 2022). Attempts to understand when and why label smoothing is effective can be traced back to Pereyra et al. (2017) and Müller et al. (2020), which respectively relate label smoothing to entropy regularization and show that it can lead to more closely clustered learned representations. Müller et al. (2020) also show that label smoothing can improve calibration but hurt distillation performance; Zhu et al. (2023); Xia et al. (2024) further show that these improvements in calibration do not translate to better selective classification.
On the theoretical side, Lukasik et al. (2020) study the relationship between label smoothing and loss correction techniques used to handle label noise, and show that label smoothing can be effective for mitigating label noise. Liu (2021) extends this line of work by analyzing how label smoothing can outperform loss correction in the context of memorization of noisy labels. Wei et al. (2022) further show that negative label smoothing can outperform traditional label smoothing in the presence of label noise. Xu et al. (2020) provide an alternative theoretical perspective, showing that label smoothing can improve convergence of stochastic gradient descent.
@@
On the other hand, Guo et al. (2019) and Chidambaram et al. (2021) show that the regularization terms introduced by the Mixup loss can also have a negative impact due to the fact that Mixup-augmented points may coincide with existing training data points, leading to models that fail to minimize the original risk. Mixup has also been studied from the perspective of calibration (Thulasidasan et al., 2019), with theoretical results (Zhang et al., 2021; Chidambaram & Ge, 2023) showing that Mixup training can prevent miscalibration issues arising from ERM.
Recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM. Most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al. (2023) and Zou et al. (2023) both showing that Mixup training can learn multiple features in a generative, multi-view (Allen-Zhu & Li, 2021) data model despite ERM failing to do so.
Neural Collapse. Papyan et al. (2020) showed that the last layer activations of trained neural networks collapse to their class means, in such a way that they are maximally separable. Kornblith et al. (2021); Zhou et al. (2022); Guo et al. (2024) all investigate the interplay between label smoothing and this effect, with Kornblith et al. (2021) showing that label smoothing can lead to more separable last layer representations (although worse linear transfer performance) and Zhou et al. (2022); Guo et al. (2024) showing that it can also lead to faster convergence to these collapsed representations. We corroborate these results with our experiments in Section 4.1, where we show that both Mixup and label smoothing lead to much lower total variance in last layer activations compared to standard cross-entropy.
@@ 2 Preliminaries
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(X_{1:2},Y_{1:2})\sim\pi\otimes\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h(\lambda,g,X_{1:2},Y_{1:2})\right].$ (2.8)
@@ 3.1 Linear Binary Classification
We begin first with the binary setting, in which we will consider a data model where a subset of the input dimensions correspond to a low variance feature and the complementary dimensions correspond to a high variance feature that is more separable. We emphasize that our notion of “feature” here does not correspond to explicit feature vectors that are fixed per class like in prior work, we simply designate subsets of the dimensions as features for simplicity. This data model can be interpreted in multiple ways: depending on the context, we may wish to learn either both of the features present in the data or simply hone in on the low variance feature (for example, identifying a stop sign with many different backgrounds).
Although this setting seems simplistic – one may object that the variance differences in the input can be handled with suitable normalization/whitening – the insights from our data model can be applied to learned features (i.e. intermediate representations in a deep learning model) where such modifications are less straightforward, as we demonstrate in the experiments of Section 4.
Our main results in this section show that doing any kind of label augmentation (label smoothing or Mixup) in our low variance/high variance feature setup will lead to a model that has only learned the low variance feature, whereas minimizing the binary cross-entropy with non-zero weight decay requires learning the high variance feature due to its greater separability.
@@
For convenience, for a vector $v\in\mathbb{R}^{d}$, we will use $v\in\mathcal{L}$ to mean that only the $\mathcal{L}$ dimensions of $v$ are non-zero. Definition 3.1 only treats $\mathcal{L}$ and $\mathcal{H}$ as placeholders; this is because our weight decay result is insensitive to variance assumptions and only depends on differences in the separability of the dimensions $\mathcal{L}$ and $\mathcal{H}$, as we indicate below.
@@
Assumption 3.2 is strong, but it actually does not imply linear separability, only that the $\mathcal{H}$ dimensions are in a sense better than the $\mathcal{L}$ dimensions for classification. We explore a simple 2-D distribution illustrating Assumption 3.2 in Appendix C.1.
@@
On the other hand, we will now show that once we introduce variance assumptions on $\mathcal{L}$ and $\mathcal{H}$, this phenomenon does not occur when minimizing the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$. In both cases, the optimal solutions have arbitrarily small correlation with the dimensions in $\mathcal{H}$ as the distribution of $YX$ concentrates. [^1] In order for the variance of $YX$ to go to 00, we need the classes to be balanced; however, we can change our assumptions to be in terms of the conditional variance of $X$ to remove this requirement. For these results, we require the following assumptions (but no longer need Assumption 3.2).
@@
We assume that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\Sigma_{YX,\mathcal{H}}\succ\rho\mathrm{I}_{d}$ for some $\rho>0$. Here $\Sigma_{YX,\mathcal{H}}$ denotes the covariance matrix of $\Sigma_{YX}$ restricted to those rows and columns in $\mathcal{H}$.
The first part of Assumption 3.5 just ensures it is possible to obtain a good solution using the dimensions in $\mathcal{L}$ while the second part codifies the idea of $\mathcal{H}$ being a high (at least non-zero) variance feature. Observe that we have made no direct separability assumptions on the data, although it is true that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ the class-conditional supports are guaranteed to be linearly separable in $\mathcal{L}$. This means that we can consider settings where both Assumptions 3.2 and 3.5 are true, and in such settings there will be a clear separation between weight decay and label smoothing/Mixup. This kind of separation in the learned decision boundaries is visualized in Appendix C.1, which provides intuition for the next two results.
@@
Proof Sketch. Mixup differs from label smoothing in that we need to condition on $\lambda$ and then show strong convexity of the conditional loss in terms of $w^{\top}Z_{\lambda}$. We can then get a lower bound similar to the label smoothing case, but it is no longer immediate that there exists a stationary point $w\in\mathcal{L}$ that minimizes this lower bound. We prove the existence of such a stationary point by considering limiting behavior of the gradient of the loss as $w_{\mathcal{L}}$ tends to $\infty$ or $\infty$- ∞ in each component, after which the rest of the proof follows the label smoothing proof.
@@ 3.2 General Multi-Class Classification
$\ell_{\mathrm{LS},\alpha}(g)\geq\mathrm{OPT}_{\mathrm{LS},\alpha}+C\sum_{y=1}^{k}\pi_{Y}(y)\mathrm{Var}\left(\mathbb{E}[g(X)\mid Y=y]\right).$ (3.1)
@@
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)\geq\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}+C\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda}\left[\mathrm{Var}\left(\mathbb{E}[g(Z_{\lambda})\mid y_{1},y_{2},\lambda]\right)\right].$ (3.2)
@@
Additionally, our results can be viewed as generalizing the observations made by Chidambaram et al. (2023), which were that Mixup can learn multiple features in the data when doing so decreases the variance of the learned predictor. In our case, we directly show that Mixup will have much larger correlation with the low variance feature in the data as opposed to the high variance feature. Our results also do not contradict the observations of Zou et al. (2023), which were that Mixup can learn both a “common” feature and a “rare” feature in the data; in their setup the common/rare features are fixed (zero variance) per class and concatenated with high variance noise. We provide a more detailed review of the settings of Chidambaram et al. (2023) and Zou et al. (2023) in Appendix B.
@@ 4 Experiments
We now address the practical ramifications of our theoretical results. In Section 4.1, we show that the intermediate representations learned by deep learning models trained with label smoothing and Mixup do indeed exhibit significantly lower variance when compared to those learned using just weight decay, and that these lower variance representations correlate with better performance. We also show, however, that this minimum variance feature learning can be a detriment by analyzing spurious correlations in the training data in Section 4.2. We also include synthetic experiments directly verifying our theoretical setting in Appendix C.2.
All experiments in this section were conducted on a single A5000 GPU using PyTorch (Paszke et al., 2019) for model implementation. All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
4.1 Learned Low Variance Features in Image Classification
(a) CIFAR-10 Test Error
(b) CIFAR-10 Activation Variance
(c) CIFAR-10 Output Variance
(d) CIFAR-100 Test Error
(e) CIFAR-100 Activation Variance
(f) CIFAR-100 Output Variance
Figure 1: ResNet-18 final test errors, penultimate layer activation variances, and output probability variances on CIFAR-10 and CIFAR-100. Activation variance results are shown starting at epoch 25 as early epochs have larger scale oscillations in the computed variance.
We first consider image classification on the standard benchmarks of CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) using ResNets (He et al., 2015); we show results for ResNet-18 in this section and relegate further experiments for deeper architectures to Appendix C.3 (they follow the same trends). We compare the performance of training using just ERM + weight decay to that of ERM + weight decay combined with label smoothing or Mixup; final test error performance is shown in Figure 1 (a) and (d).
Due to compute constraints, we focus on known well-performing settings (Zhang et al., 2018; Müller et al., 2020) for weight decay, label smoothing, and Mixup. Namely, we take the weight decay parameter to be $5\times 10^{-4}$, the label smoothing $\alpha$ parameter to be $0.1$, and the mixing distribution for Mixup to be $\mathrm{Beta}(1,1)$ (uniform distribution). We train all models for 200 epochs using a batch size of 1024, a fixed learning rate of $10^{-3}$, and AdamW for optimization (Loshchilov & Hutter, 2019). We preprocess the training data to have zero mean and unit variance along each channel, and also include random crop and flip augmentations as is standard practice for achieving good performance.
Figure 1 shows that adding in label smoothing and Mixup to the baseline of ERM + weight decay leads to noticeable performance improvements, which corroborates existing results in the literature. The novel aspect of our results is shown in Figure 1 (b), (c), (d), and (e), where we track the mean total variance of the penultimate layer activations and output probabilities of each model on the test data over the course of training. The variance values are computed by first computing the total variance of the activations/probabilities (i.e. the sum of the variance in each dimension) for each class, and then averaging over all classes. Final test error and variance values are provided in Table 1 and 2 in Appendix C.3.1.
Essentially, we measure the average spread of activations and probabilities across classes, which is similar in spirit to what was done by Müller et al. (2020), although we directly look at variances across all classes whereas they look at projected clusters of a few specific classes. The most telling results are the model activation variances, since it is to an extent expected that per-class variance of model outputs should decrease with improvements in test error (this also corresponds to improved model calibration, which is a known consequence of training with label smoothing and Mixup). That being said, we further analyze model output variance in Appendix C.4 and show that label smoothing and Mixup decrease total output variance by a larger extent than what can be explained by changes in the target class prediction variance alone.
Overall, our results show that adding either label smoothing or Mixup to the baseline of just weight decay leads to significant decreases in the activation variances, adding credence to the idea that both methods learn low variance features. We note, however, that our results only establish a correlation between this low variance property and better test performance – it would require a significantly more in-depth empirical study to assess a causal relationship between this kind of feature learning and generalization, which we think would be a fruitful avenue for future work.
4.2 Low Variance Spurious Correlations
We now demonstrate that honing in on low variance features can also be harmful to performance via binary classification and multi-class classification tasks in which the training data is modified to have spurious correlations with the target. The introduced spurious correlations are much lower magnitude than the rest of the data, and in that sense they intuitively satisfy both Assumptions 3.2 and 3.5.
4.2.1 Binary Classification With Perturbed Training Data